Vectors

A football player runs directly down field for 35m before turning to the right at an angle of 25 from his original direction and running a additional 15m beforegetting tackled.What is the magnitude and directoin of the runners total displacement
Answer #1
initially the foot ball player travells a distance of 35m alonthe Y axis so we can writeinitial dispalcement as 35mhe now turns at an angle 250 from his initialdisplacement &travells an additional distance of 15m,so thefinal displacement is15msin250+15mcos250so the total displacement is = 35m+15msin250+ 15mcos250= 6.34m+ 48.6mthus the magnitude of resultant displacement is ||=√(6.34m)2+(48.6m)2=49mthe direction is given by θ = tan-1(6.34/48.6)=7.430thus therunner final displacement is7.430with the initial displacement.
answered by: de de